May 28, 2013 Solenoid Meeting
Tuesday May 28,2013 at 11am At Jefferson Lab, the meeting will be held in room ARC 231
- Discussion on Quench Details
- Plans for low current testing
- Work accomplished since Quench
- Power Supply ramp smoothing program
- Power Supply resolution hardware fix (0-1.5 A/sec)
- Voltage tap checks/fixes
- Short to ground location
- Power Supply hose replacement
- Plans for low current testing
- 15A Commissioning Plan Media:15A Plan.pdf
- 140A Commissioning plan
- Cool Down Plans
- Path Forward .......
- Documents available at M:\halld_engineering\Solenoid\Commissioning\Response to Quench
Probir and Renuka discussed/presented their findings on the quench analysis and comparison to the LASS configuration. Based on this analysis there should have been a minimum of 1K margin on the superconductors for all coils. Bob Flora suggested that a momentary turn to turn short may be able to cause the quench. Eugene will calculate this affect. After much discussion there was speculation that the superconductors have local areas that are not as good as the short samples tested at SLAC. Still no consensus on the cause of the quench. The refrigerator has carbon and oil issues that Jonathan is working through. Jonathan is doubtful that we will get much time from the refrigerator without major work (~2 months).
Our Path forward from the discussion is as follows;
- 1. Perform 15A tests with power supply connected to the magnet as outlined in commissioning plan D00000-04-02-P002 Rev B
- 2. George to update commissioning plan to rev C for 140A testing
- 3. Jonathan to continue to work on the refrigerator to get the magnet to 4.5K
- 4. Perform 140A testing when plan is approved and magnet cold
In the mean time, the following will be done in parallel to prepare for high current testing;
- 1. Eugene to calculate if a momentary turn to turn short could have caused a quench
- 2. George and Probir to determine whether a quench can be captured and the consequences to the magnet if only 1 coil quenches and whether or not a change in dump resistors would be helpful. This would take into consideration the maximum safe voltage for the magnet.
- 3.Determine a safe current to run the magnet at that will also satisfy physics requirements and DOE milestones. 1350A appears to satisfy the physics and DOE requirements. We need to come up with the determination of whether or not this gives us a safe operating margin. Possibly a 1000 liter Dewar needed at all times during the high current run.
- 4. Modify the power supply to tune for a 25H magnet. Danfysik is to supply the hardware and instructions on how to do so.
- 5. George to prepare a procedure to manually adjust the quench detector as the current is increased.
- 6.Eugene/George to prepare a final write up on the quench and it's probable or speculated causes.
Where do we go from here?
Taking into consideration all the problems with the refrigerator and the overall Hall D schedule (and re-baseline schedule), it would be prudent to move forward on all testing and magnetic field mapping as soon as possible in a safe manner for the magnet. In order to do this, all of the low current testing would have to show that the magnet is still viable, and the items listed above for high current testing need to be completed. Based on this, the following is my suggestion;
- 1. Complete 15A testing
- 2. Cool the magnet to 4.5K
- 3. Test to 140A
- 4. Ramp to 1400A and hold for 30 min (with Dewar attached)
- 5. Ramp down to 1350A and map the magnetic field. (with Dewar attached)
- 6. Ramp down and back up to 1350A (with Dewar attached)
- 7. Commissioning complete - warm magnet and make scheduled repairs to the magnet and refrigerator.
For future production running of the magnet, we will need to develop the quench detection system so it will not need manual intervention for adjusting for inductance imbalance as the current ramps up. We will work with Bob Flora on this. Also, the refrigerator will need to be replaced or a major overhaul will need to be done. This decision will need to be made after initial testing.